Fermi National Laboratory
 
Operation Summary for December 14, 2011 - December 16, 2011
 

Wednesday December 14

The day shift began with Operations monitoring the Linac, Booster, and Main Injector (MI), and with the NuMI and MiniBooNE experiments taking beam.

During the day shift, Booster personnel began studies at 9:35 AM. Operations established beam to Meson FTBF experiment T-1015 around 10:40 AM. Operators rebooted a NuMI BPM VME crate for a NuMI BPM (NBPM65) at 10:41 AM.

During the evening shift, a Linac person began a 10-minute 750 KeV Line study at 5:32 PM. Operators rebooted a NuMI BPM (NBPM6N) at 6:20 PM. At 11:04 PM, the I- Source’s output began dropping; Operators increased the arc supply voltage. Operators rebooted NBPM65 and NBPM6N at 11:23 PM.

Thursday December 15

The midnight shift began with Operations sending beam to the NuMI and MiniBooNE.

During the midnight shift, at 1:10 AM, an I- Source vacuum pump tripped off; it reset. At 2:42 AM, the NuMI beam permit tripped off for the first of two times in about four hours due to a MI RF gap indication; it reset.

During the day shift, at 8:07 AM, the MI gap clearing kickers tripped off: one with a Thyratron pre-fire indication and the other with a timing and sequence error; everything reset. A Booster person began a study at 8:20 AM. At 8:59 AM, the Booster Correction magnet Element Package, B:CEP400, tripped off; it reset. Operations established beam to T-1015 at 10:21 AM.

During the evening shift, Operators rebooted NBPM6N at 5:09 PM. At 10:32 PM, the I- Source’s output began dropping; Operators increased the arc supply voltage. Operations disabled T-1015 beam at 10:41 PM. At 11:36 PM, the I- Source’s output began dropping; Operators increased the arc supply voltage again. At 11:57 PM, operators rebooted NBPM65.

Friday December 16

The midnight shift began with Operations sending beam to the NuMI and MiniBooNE.

During the midnight shift, Operators completed a MI quadrupole magnet tune at 3:30 AM. Operators completed a 400 MeV Line tune, a 750 keV Line tune, and an I- Source tune at 5:58 AM. Operators completed a Booster injection tune at 6:22 AM. At 7:19 AM, a Linac quadrupole magnet power supply’s (L:QPS107) output current began drifting; Operators swapped out the control cards, but it didn’t help. A Linac person reported that the power supply had to be replaced.

Operations Coordinator (OpsCo) Plans for Today and the Weekend

The OpsCo reported that there would be a four-hour access for various jobs early next week. A NuMI person reported they had their third best week for Protons on Target.

 

Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now.

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.

last modified 12/16/2011    email Fermilab
Security, Privacy, Legal  |  Use of CookiesFermi National Accelerator Laboratory